OpenStudy (anonymous):

please help give the function by h(x) = x + 1

OpenStudy (anonymous):

And do what with it?

OpenStudy (anonymous):

sorry h(0) = h(20) = h(-4)= h(6) = h(x+13) =

OpenStudy (anonymous):

Quantum Torch you still there